show steps pleaseeeee! ~(^.^)~
6x-5z=17
5x-y+3z=-1
2x+y=-41
find x,y,and z

Answer by ankor@dixie-net.com(22740)

6x - 5z = 17
5x - y + 3z =-1
2x + y = -41
:
This can be solved with elimination, rewrite the equation and add
6x + 0y - 5z = 17
5x - y + 3z = -1
2x + y + 0z = -41
------------------adding eliminates y
13x - 2z = -25
:
Multiply this equation by 5, and the 1st equation by 2
65x - 10z = -125
12x - 10z = 34
-----------------subtraction eliminate z, find x
53x = -159
x = -159/53
x = -3
:
Find y using the 3rd equation
2(-3) + y = -41
-6 + y = -41
y = -41 + 6
y = -35
:
Find z using the 1st equation
6(-3) - 5z = 17
-18 - 5z = 17
-5z = 17 + 18
-5z = 35
z = 35/-5
z = -7
:
x=-3, y=-35, z=-7
:
:
Confirm this in the 2nd equation (5x - y + 3z =-1)
5(-3) - (-35) + 3(-7) =-1
-15 + 35 - 21 = -1, equality reigns!